OpenStudy (anonymous):

Your piggy bank has a total of 43 coins in it; some are dimes and some are nickels. If you have a total of $3.60, how many dimes do you have?

OpenStudy (mathhelp346):

so you know you have 43 coins. An equation for that would be n+d=43

OpenStudy (lgbasallote):

you take this one @Mathhelp346 ^_^

OpenStudy (mathhelp346):

nvm the second equation should be 10d+5n=160

OpenStudy (anonymous):

26 x 17 =3.60 im close... 3.45

OpenStudy (mathhelp346):

then you would solve for either d or n in the first equation. i'm going to do n. you would subtract d from both sides, leaving you with n=43-d

OpenStudy (mathhelp346):

you would sub n into the second equation and solve

OpenStudy (mathhelp346):

do you know what to do from here?

OpenStudy (anonymous):

29 dimes and 14 nickles.

OpenStudy (mathhelp346):

i don't think that's right...

OpenStudy (anonymous):

right?

OpenStudy (mathhelp346):

that would add up to 360. you need 160

OpenStudy (anonymous):

no the question is asking $3.60

OpenStudy (mathhelp346):

oh my bad. your answer is right.